# GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering

### v1

@@ Abstract
Pre-trained Language Models (PLMs) have shown remarkable performances in recent years, setting a new paradigm for NLP research and industry. The legal domain has received some attention from the NLP community partly due to its textual nature. Some tasks from this domain are represented by question-answering (QA) tasks. This work explores the legal domain Multiple-Choice QA (MCQA) for a low-resource language. The contribution of this work is multi-fold. We first introduce JuRO, the first openly available Romanian legal MCQA dataset, comprising three different examinations and a number of 10,836 total questions. Along with this dataset, we introduce CROL, an organized corpus of laws that has a total of 93 distinct documents with their modifications from 763 time spans, that we leveraged in this work for Information Retrieval (IR) techniques. Moreover, we are the first to propose Law-RoG, a Knowledge Graph (KG) for the Romanian language, and this KG is derived from the aforementioned corpus. Lastly, we propose a novel approach for MCQA, Graph Retrieval Augmented by Facts (GRAF), which achieves competitive results with generally accepted SOTA methods and even exceeds them in most settings.
GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering
Cristian-George Craciun1,3thanks: cristian.craciun@tum.de Răzvan-Alexandru Smădu3, Dumitru-Clementin Cercel3thanks: dumitru.cercel@upb.ro, Mihaela-Claudia Cercel2,4 1Technical University of Munich, Munich, Germany 2Paris 1 Panthéon-Sorbonne University, Paris, France 3National University of Science and Technology POLITEHNICA Bucharest, Faculty of Automatic Control and Computers, Bucharest, Romania 4University of Bucharest, Bucharest, Romania
@@ 1 Introduction
With the advent of the Transformer architecture (vaswani2017attention) and its parallelization capabilities, people have sought to scale the models in terms of parameters (kaplan2020scaling), unlocking new abilities for machine learning models to perceive natural language at a finer level. Since then, Large Language Models (LLMs) have been pre-trained on various tasks, achieving better performances on downstream tasks (gururangan2020don) than their non-trained counterparts.
QA represents a family of NLP downstream tasks explored in various works (zhong2020jec; rogers2023qa; singhal2023towards). Some QA tasks can be formulated as open-ended questions that require an elaborate answer, which may or may not include the rationale behind it. Other works focus on retrieving the answer from a given context, in which case the answer can either be ad litteram found in the provided context (zaib2021bert) and a model has to identify the location at which the answer occurs, or the model has to infer the correct answer based on an understanding or interpretation of the context at hand, task which bears the name Machine Reading Comprehension (MRC) (baradaran2022survey).
The aforementioned tasks have been explored for various domains and languages (askari2022expert; louis2021statutory; hoppe2021towards). This work focuses on MCQA for the Romanian legal domain. Legal QA represents a very promising area of research due to the insights it can provide in tackling various other problems, as well as, from a practical standpoint, by benefiting the average citizen of a nation with the possibility of finding the answer to a legal inquiry together with experts in the field who will achieve increased productivity.
Additionally, available natural language data are scarce for the Romanian language. The only MCQA work was performed in (dima2024roqllama), which proposed a Romanian medical MCQA dataset. Although the legal domain for the Romanian language has been tackled in some works in the past (masala2021jurbert; masala-etal-2024-improving), there is no open-source datasets available. This motivated us to introduce new comprehensive resources for the Romanian language.
To summarize, the contributions of this paper are the following:
  • We release JuRO [^1] https://github.com/craciuncg/JuRO, the first open-source legal dataset available for the MCQA Romanian.
@@
  • We release Law-RoGfootnotemark: , the first KG for the Romanian language, a KG for the legal field.
  • We present a novel algorithm for MCQA called GRAFfootnotemark: .
  • We provide a comprehensive evaluation of our proposed MCQA dataset using SOTA methods and our proposed method.
@@ 2 Related Work
A natural follow-up for the QA domain represented the adaptation of this research branch to downstream fields that require specific knowledge, particularly examinations. Also, the plethora of freely and publicly available legal texts made legal question answering a convenient start for testing diverse engineering solutions and targeted algorithms, allowing for the development of proof-of-concept methods.
2.1 Legal QA
English Work. There have been initial traditional approaches to legal question answering for bar exam questions (john2017solving) using BiLSTM networks. The remarkable achievements of the Transformer models (vaswani2017attention) motivated people to use them in this domain as well. One work revolved around privacy-related law shankar2023privacyglue, using various sources for building their benchmark dataset, besides using several legal BERT (devlin2018bert) models for their experiments. Another work has focused on IR of legal paragraphs, designing a new approach using a multi-stage pipeline, incorporating the BM25 algorithm (robertson1976relevance) and BERT for semantic similarity.
Multilingual Work. Other languages also took the initiative, and linguistic research groups attempted to solve legal tasks in their languages using diverse approaches. One work in the Chinese language addressed the legal field, featuring an MCQA dataset along with a comprehensive evaluation of diverse methods zhong2020jec. An Italian research group built a new dataset for international private law sovrano2021dataset. A Vietnamese research team (kien-etal-2020-answering) has proposed a legal QA system as well as an annotated corpus for testing their solution.
Romanian Work. For the Romanian language, one particularly useful work is performed in (masala2021jurbert), which proposes jurBERT, a BERT model adapted to the Romanian legal domain through an extensively large corpus of legal text, which amounts to 160GB, according to the authors. They also tested the model on downstream tasks, in which the smaller jurBERT model surprisingly achieved better results. Following this result, the authors attempted to build a judicial prediction system (masala-etal-2024-improving) using the above-mentioned model.
@@ 2.1 Legal QA
Table 1: General statistics for CROL and KG.
2.2 Graph Based Approaches
The literature features variable approaches when it comes to KG-based QA. We identify 2 main classes of approaches: symbolic and numeric. Since the KG is a symbolic structured representation of factual knowledge, some works revolved around symbolic methods. One work (chakraborty2024multi) explored multi-hop QA approaches using an LLM-KG-based algorithm, which identified the correct answer. The numeric approaches attempt to make use of learned numeric representations of the data to make predictions. One such work is performed in (he-etal-2022-dialmed), which proposes a medical dialogue dataset as well as a method for using a medical graph (wang2019kgat) for predicting the corresponding medication given patient-doctor dialogue. Our work attempts to create a synergy between the two approaches, as well as propose the datasets and the KG for future uses.
3 Novel Resources
Dataset Size Category Language
PrivacyQA (ravichander2019question) 1,750 Span English
PolicyQA (ahmadetal2020policyqa) 714 Span English
JEC-QA (zhong2020jec) 26,367 Multi-Choice Chinese
COLIEE-21 (rabelo2022overview) 887 Binary Japanese
BSARD (louis2021statutory) 1,100 Article Retrieval French
[…]
Table 2: General statistics for CROL and KG.
JuRO. We introduce a new dataset for legal QA, which encompasses 3 different types of examinations: Entrance into the judicial system, entrance into the Bar, and Promotion exams for judicial positions. We are the first to propose such a dataset for the Romanian language and release a Romanian openly available legal dataset. We hope this will open opportunities for the Romanian, multilingual, and low-resource language research. More details for this dataset as well as other resources can be found in Appendix LABEL:appendix:a.
@@ 3 Novel Resources
CROL. Corpus for Romanian Law represents a collection of gathered legal documents for all of the law branches mentioned in Figure LABEL:fig:juropie. It can mainly serve as a knowledge base for IR techniques for Romanian legal tasks. There has been past work on creating a Romanian legal corpus, such as the Marcell project (varadi2020marcell), which aimed at creating a multilingual legal corpus, including Romanian law. However, this represents an annotated text corpus useful for NER training in a legal context as well as pertaining, but did not make a distinction between "documents that are in effect and those that are not.” We have made a clear separation between legal documents and their updates.
@@
Law-RoG. We present the first-ever KG for the Romanian language. This KG is built upon the CROL corpus via entity-relation extraction. Particularly, following this work (edge2024local), we use a single prompt for identifying named entities as well as the relations between them and instruct a LLM to output entity-relation-entity triplets in our desired format using in-context (brown2020fewshot) learning abilities that LLMs exhibit via few-shot prompting. We opted for this approach since the Romanian NLP lacks resources for building a specialized pipeline for entity and relation extraction, especially for the legal domain. To make sure this model would generally produce factually and coherently correct information, we sampled 10 documents and their corresponding generated relations from each legal domain, and used a total of 5 human NLP experts, and we all reached a consensus that the outputs were coherent enough and did not hallucinate beyond the given document, and even though not perfect, we concluded they are more than enough for almost all related NLP applications.
3.1 Data Collection
JuRO dataset represented an extensive work for it to reach the final current form. The data was found in multiple forms and had to resort to multiple tools and stages. Where possible, we used web scrapping to extract questions and their corresponding answers. Further, we scrapped PDF documents with textual information available. Lastly, we used OCR to extract the underlying text from the scanned PDFs. In addition, we went for manual verification of data because some scans presented high-level damages of noise, which made OCR impractical, and even though manual extraction was tedious, it was more time economical than going through the OCR process at this stage. All exams are publicly available on the official portals of the institutions responsible for examinations.
The CROL corpus was extracted exclusively via scrapping from the official Ministry of Justice portal, department of the Romanian Government, in a different manner as opposed to the Marcell project (varadi2020marcell), by ensuring an organized crawling of the portal and extracting legal documents along with their distinct legal articles and their updates.
@@ 4 GRAF
We introduce a novel approach for retrieving information from a KG, which we applied to legal MCQA. The same principles discussed regarding claim checking and validation can be applied to other tasks requiring factual knowledge.
@@ 4.1 Problem Formulation
At first glance, we are presented with questions that may have only 1 correct answer for a dataset, and in other scenarios, a question can have up to 2 correct choices. We will make a distinction between these 2 settings and will formulate the goal of the problem according to the architecture of the proposed model. In any case, the dataset can be represented as $\mathcal{D}=\{e_{i}=(Q_{i},C_{i},T_{i})|i\in\{1..|\mathcal{D}|\}\}$. Each triplet entry $e_{i}$ contains the question body $Q_{i}$ in textual form, the set of possible answer choices $C_{i}=\{C^{1}_{i},C^{2}_{i},C^{3}_{i}\}$ and the set of target answers $T_{i}$ with $|T_{i}|\leq 2$ and $T_{i}\subset C_{i}$. Each answer choice is a tuple $C^{j}_{i}=(\sigma^{j}_{i},\epsilon^{j}_{i})$ where $\sigma^{j}_{i}$ is the symbol associated with the text content $\epsilon^{j}_{i}$ of the answer choice. We will now look at 2 classes of models we also experimented with and formulate the learning goal.
where $j>k$ and $t^{0}_{i}$ represents the empty sequence. The $\mathcal{Z}$ function maps a given triplet to a sequence that can be processed by the given probabilistic model, known as the model prompt. The prompt has the purpose of facilitating and guiding the model towards a lower on-average negative log-likelihood and consequently, the correct answer. Even though this work also explores sequence-to-sequence models, the ones we chose in particular also generate the output in an autoregressive fashion via the decoder model, and thus, our previous discussion still holds.
Encoder Models. These models showed excellent performances on classification tasks despite their relatively smaller size in practice. They feature a good semantic understanding of a given sequence via their pretraining objectives. For instance, BERT (devlin2018bert) featured word level and sentence level pretraining, which allowed it to gain a semantic understanding of language at these levels. However, they do not exhibit symbol level correlation (robinson2023leveraging) as LLMs do, and thus, we resort to using their semantic understanding of textual sequences for outputting a number that represents the degree to which a given choice is correct given a question. We consider two learning goals for these models, the binary cross entropy minimization for models outputting probabilities:
$\mathcal{L}_{1}=-(o^{j}_{i}\log(y^{j}_{i})+(1-o^{j}_{i})\log(1-y^{j}_{i}))$ (2)
where $o^{j}_{i}=\mathbbm{1}_{T_{i}}(C^{j}_{i})$ and $y^{j}_{i}$ is the model output probability. We also use the cosine similarity embedding loss to align a given question with the correct answer choice:
$\mathcal{L}_{2}=(1+o^{j}_{i})(1-y^{j}_{i})+(1-o^{j}_{i})y^{j}_{i}$ (3)
For the inference, we consider the question along with the set of choices and select the top $|T_{i}|$ scores in the following way:
@@
$Y_{i}^{*}=\text{TopK}(Y_{i},|T_{i}|)$ (4)
Figure 2: Translated sample question from the JuRO dataset and a showcase of GRAF on this sample. In this case, the entities "Self-defense" and "Unauthorized Access" have the best alignment with the entities in choice "C" thus it is most likely to be the correct answer.
Data: $Q$ Question, $C$ Choices
$G$ Knowledge Graph
Result: $P$ Choice Proababilities
$P\leftarrow\text{[]}$]
$C\leftarrow$ split_choices($C$)
for $c_{i}\in C$ do
        /* Query Claim Extraction model and obtain $CG$ Claim Graph */
       $CG\leftarrow\text{claim\_extract}(Q||c_{i})$
        /* Sample $G$ into Sub Graph $SG$/
       $SG\leftarrow\text{sample\_graph}(Q||c_{i})$
        /* Encode $SG$ and $CG$/
       ${}{}SGE,CGE\leftarrow\text{\{\}},\text{\{\}}$} , {}
       for $(E_{a},r_{ab},E_{b})\in SG$ do
             $SGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
       end for
      for $(E_{a},r_{ab},E_{b})\in CG$ do
             $CGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
       end for
      $SGE,CGE\leftarrow\text{GAT}(SGE),\text{GAT}(CGE)$
       for $(h^{i}_{c},h^{i})\in CGE\times SGE$ do
             $R^{ij}\leftarrow\cos(h^{i}_{c},h_{i})$
       end for
      $\bar{H}\leftarrow Rh$
       $c\leftarrow\text{enc}(Q||C)$
       $c_{\text{final}},H_{\text{final}}\leftarrow\text{SelfAttention}(c||\bar{H})$
       $P\text{.append}(c_{\text{final}})$
end for

### v2

# GRAF: Graph Retrieval Augmented by Facts for Romanian Legal Multi-Choice Question Answering

@@ Abstract
Pre-trained language models have shown remarkable performances in recent years, setting a new paradigm for natural language processing (NLP) research. The legal domain has received some attention from the NLP community partly due to its textual nature. Question-answering (QA) systems represent some of the tasks in this domain. This work explores the legal Multiple-Choice QA (MCQA) for a low-resource language, Romanian. The contribution of this work is multi-fold. We introduce JuRO, the first openly available Romanian legal MCQA dataset, comprising 10,836 questions from three examinations. Along with this dataset, we introduce CROL, an organized corpus of laws with a total of 93 distinct documents with their modifications over 763 time spans, which we used for information retrieval techniques in this work. Additionally, we construct Law-RoG, the first graph of legal knowledge for the Romanian language derived from the corpus mentioned above. Lastly, we propose a novel approach for MCQA, namely Graph Retrieval Augmented by Facts (GRAF), which achieves competitive results with generally accepted state-of-the-art methods and even exceeds them in most settings.
GRAF: Graph Retrieval Augmented by Facts for Romanian Legal Multi-Choice Question Answering
Cristian-George Craciun1,2thanks: cristian.craciun@tum.de Răzvan-Alexandru Smădu1, Dumitru-Clementin Cercel1thanks: dumitru.cercel@upb.ro, Mihaela-Claudia Cercel3,4 1National University of Science and Technology POLITEHNICA Bucharest, Faculty of Automatic Control and Computers, Bucharest, Romania 2Technical University of Munich, Munich, Germany 3Paris 1 Panthéon-Sorbonne University, Paris, France 4University of Bucharest, Bucharest, Romania
footnotetext: Under review.
@@ 1 Introduction
Question-answering (QA) represents a family of downstream natural language processing (NLP) tasks explored in various settings (Zhong et al., 2020; Rogers et al., 2023; Singhal et al., 2023). Some QA tasks can be formulated as open-ended questions that require an elaborate answer, which may or may not include the rationale behind it (Chen et al., 2017; Labrak et al., 2022). Other studies focus on retrieving the answer from a given context. In this scenario, the answer can either be found explicit in the provided context, and a model has to identify the location at which the answer occurs (Zaib et al., 2021) or infer the correct answer based on understanding and interpretation of the context at hand, a task known as machine reading comprehension (Baradaran et al., 2022).
The tasks mentioned above have been explored for various domains and languages (Louis and Spanakis, 2022; Hoppe et al., 2021; Askari et al., 2022; Sen et al., 2022; Ekram et al., 2022). Our work focuses on MCQA for the Romanian legal domain. Legal QA represents an emerging area of research due to the insights it can provide in tackling various other problems, such as the dialogue system He et al. (2024), as well as, from a practical perspective, by benefiting the average citizen of a nation with the possibility of finding the answer to a legal inquiry together with experts in the field who will achieve increased productivity.
Additionally, available natural language corpora are scarce for the Romanian language. The only MCQA work was performed by Dima et al. (2024), which proposed a Romanian medical MCQA dataset. Although the Romanian legal domain has been tackled in the past (Masala et al., 2021, 2024), no open-source datasets are available. This motivated us to introduce new comprehensive resources for the Romanian language.
In summary, the contributions of this paper are the following:
  • We release JuROhttps://github.com/craciuncg/JuRO, the first open-source legal dataset available for the Romanian MCQA.
@@
  • We release Law-RoGfootnotemark: , the first knowledge graph for the Romanian legal domain.
  • We present a novel algorithm for the MCQA task called GRAFfootnotemark: .
  • We provide a comprehensive evaluation of our proposed MCQA dataset using state-of-the-art methods and our proposed method.
@@ 2 Related Work
2.1 Legal-Domain Question Answering
English QA. Initially, traditional legal QA systems were based on information retrieval techniques, rule-based systems, support vector machines Kim et al. (2014, 2017) and shallow convolutional networks (CNNs) Kim et al. (2015b) to answer legal questions from bar exam questions (Kim et al., 2015a). Later, deep neural networks based on CNNs Do et al. (2017) and BiLSTM John et al. (2017) were used with improved performance. The remarkable achievements of the Transformer models (Vaswani et al., 2017) motivated the use in this domain. Shankar et al. (2023) revolved around privacy-related law using various sources to build their benchmark dataset and evaluated several legal BERT (Devlin et al., 2019) models. Similarly, Hendrycks et al. (2021b) evaluated multiple BERT-based architectures on a newly proposed dataset, showing that the volume of training data and the architecture influence the performance. To support these findings, new datasets were proposed Sovrano et al. (2021), including multidisciplinary large datasets Hendrycks et al. (2021a).
Other Languages. Other languages also took the initiative in the early days. For example, the COLIEE 2015 shared task Kim et al. (2015a) proposed the first legal QA task on the Japanese legal documents provided in both Japanese and English. Bach et al. (2017) proposed an analysis of the Vietnamese transportation laws and evaluated a CRF-based system as a prerequisite stage for QA. An encoder-decoder architecture was proposed by Kien et al. (2020), containing word embeddings, convolution, and attention layers. The model was evaluated on a dataset containing 6,000 Vietnamese legal questions, outperforming existing retrieval-based methods. Later, Vuong et al. (2023) proposed an end-to-end retrieval-based system that used a pre-trained BERT model on weakly labeled data. Zhong et al. (2020) addressed the Chinese legal field, featuring an MCQA dataset containing practice exercise questions, and a knowledge database. The comprehensive evaluation of various methods, such as transformers, attention, and distant supervision, showed a significant gap until the human-level performance was achieved.
2.2 Romanian Legal Domain
[…]
2.3 Knowledge Graph-Based Question Answering
The literature features variable approaches when it comes to KG-based QA. We identify two main classes of approaches: symbolic and numeric. Since KG is a symbolic structured representation of factual knowledge, some works revolved around symbolic methods. Chakraborty (2024) explored multi-hop QA approaches using an LLM-KG-based algorithm, which identified the correct answer. The numeric approaches attempt to employ learned numeric representations of the data to make predictions. One such work is performed by He et al. (2022), which proposes a medical dialogue dataset as well as a method for using a medical graph (Wang et al., 2019) for predicting the corresponding medication given patient-doctor dialogue. Our work attempts to create a synergy between the two approaches and proposes new datasets and a KG for future uses.
3 Novel Resources
Dataset # Examples Category Language Public
PrivacyQA (Ravichander et al., 2019) 1,750 Span English
PolicyQA (Ahmad et al., 2020) 714 Span English
JEC-QA (Zhong et al., 2020) 26,367 Multi-Choice Chinese
COLIEE-21 (Rabelo et al., 2022) 887 Binary Japanese
BSARD (Louis and Spanakis, 2022) 1,100 Article Retrieval French
[…]
Table 1: Comparison of JuRO with other existing datasets.
Corpus Size Category Language Public
Marcell (Váradi et al., 2020) 317k documents, 774M tokens Pretraining Multilingual *
RoJur (Masala et al., 2021) 11M entries, 160GB Cases Romanian
Ostendorff et al. (2020) 200k court rulings Rulings German
Collarana et al. (2018) 62 pages, 64 sections, 24k words Span English
Kien et al. (2020) 8.5k documents, 118k articles Article Vietnamese
[…]
    BG, HR, HU, PL, RO, SK, SL
Table 2: Comparison of CROL with other datasets from various works.
3.1 JuRO
We introduce a new dataset for the legal QA, called JuRO. It contains past examination questions from all of the legal branches examined in Romania. It represents the first dataset of its kind, which we release to the public (see also Table 3).
Dataset Construction. The data is extracted from various official examination portals. Some of the subjects were extracted using OCR, and to avoid errors, we manually inspected all of the data, and thus, the data presents minimal damage. Each entry essentially consists of the body in which a theoretical question regarding a legal aspect is asked and three possible answer choices labeled A, B, and C, out of which at most two answers are correct.
Statistics. The dataset contains questions from three types of examination: entrance into the judicial system, entrance into the bar, and promotion exams for judicial positions. We present the distribution among legal domains and possible answer combinations in Figure 5 from Appendix A. Promotion exams have three possible choices with a single correct answer, whereas the others have up to two possible correct answers. The distribution among correct answers is generally balanced, with a small exception for bar exams. However, it is noteworthy that questions with a single correct answer are predominant and balanced. Data analysis for the JuRO dataset is presented in Appendix A.
JuRO vs Existing Work. We introduce a new dataset for the Romanian legal QA, which encompasses three different types of examinations: entrance, bar, and promotion. We are the first to propose a legal MCQA dataset for the Romanian language and release it to the public. We hope this will open opportunities for future research on Romanian, multilingual, and low-resource language settings. In Table 3, we compare our work against other legal datasets. Although less than half the size of the JEC-QA dataset Zhong et al. (2020), it is larger than other existing datasets for legal QA from other languages.
3.2 CROL
@@
Table 3: General statistics for CROL and Law-RoG KG.
Corpus for Romanian Law (CROL) represents a collection of gathered legal documents for all of the law branches such as civil, penal, work, administration, commercial, family, and international.
Dataset Construction. The CROL corpus construction used the official Ministry of Justice department portalhttps://legislatie.just.ro/ for crawling all of the laws from the covered branches in the JuRO dataset. All resources mentioned above have been extracted from official sources of national state institutions. Therefore, the language is formal and presents little to no grammatical errors.
Statistics. CROL represents a collection of organized legal documents, amounting to 93 distinct laws and 768 different versions of these laws. It contains 330k articles totaling about 31.5M words with a vocabulary of 78.3k words. Statistics are also presented in Table 3.2 and in Appendix A.
@@
CROL vs Existing Work. Our corpus can mainly serve as a knowledge base for information retrieval (IR) techniques for Romanian legal tasks. There has been past work on creating a Romanian legal corpus, such as the Marcell project (Váradi et al., 2020), which aimed at creating a multilingual legal corpus, including Romanian law. However, this represents an annotated text corpus useful for NER training in a legal context as well as pertaining, but does not make a distinction between "documents that are in effect and those that are not.” We have made a clear separation between legal documents and their updates. There are also other efforts Collarana et al. (2018); Kien et al. (2020); Masala et al. (2021); however, they are not available to the public (see Table 3).
3.3 Law-RoG
@@
We introduce the first knowledge graph for the Romanian language, called Law-RoG. This KG is built upon the CROL corpus via entity-relation extraction. Particularly, following the work of Edge et al. (2024), we prompt an LLM to identify named entities and the relations between them to output entity-relation-entity triplets in our desired format using in-context (Brown et al., 2020) learning abilities that LLMs exhibit via few-shot prompting (see Appendices H and I). We opted for this approach since the Romanian NLP lacks resources for building specialized pipelines for entity and relation extraction, specifically for the legal domain. To validate that the LLM produced factually and coherently correct information, we asked 5 human NLP experts to evaluate 10 randomly sampled documents and their corresponding generated relations for each legal domain. They all agreed that the outputs were coherent and did not hallucinate beyond the given document. Although not perfect, we concluded the generated relations were appropriate for almost all related NLP applications. The resulting KG spans 160k nodes and 320k edges (see Table 3.2).
@@ 4.1 Problem Formulation
The multi-choice QA problem can be formulated as follows. Consider a dataset $\mathcal{D}=\{x_{i}=(Q_{i},C_{i},T_{i})\},i=1:\left|\mathcal{D}\right|$, where each triplet entry $x_{i}$ contains the question body $Q_{i}$ in textual form, the set of $|C_{i}|$ possible answer choices $C_{i}=\{C^{k}_{i}\},{k=1:|C_{i}|}$ and the set of target answers $T_{i}$ with $|T_{i}|\leq|C_{i}|$ and $T_{i}\subseteq C_{i}$. Each answer choice is a tuple $C^{j}_{i}=(\sigma^{j}_{i},\epsilon^{j}_{i})$ where $\sigma^{j}_{i}$ is the choice label (e.g., A, B, C) of the answer, and $\epsilon^{j}_{i}$ represents its textual content. In our setting, we look at questions with $|C_{i}|=3$ choice answers and $|T_{i}|\in\{1,2\}$ correct answers, i.e., single-choice and multi-choice QA, depending on the dataset. Moreover, we investigate two classes of models designed to address these QA variations and formulate the learning goal in Appendix B.
@@
Figure 2: A showcase of how GRAF works on a sample question from the JuRO dataset translated to English. First, we construct the claim graph using an LLM that extracts the entities (nodes) and relations between them (edges). Based on the sub-KG extracted from Law-RoG and the claim graph, we determine that the entities “Self-defense” and “Unauthorized Access” have the best alignment with the entities in choice “C”, thus it is most likely to be the correct answer.
Data: $Q$ question, $C$ choices, $G$ knowledge graph
Result: $P$ choice probabilities
1 $P\leftarrow\text{[]}$]
2 for $c_{i}\in split\_choices(C)$ do
       // Query the cross claim extraction model and obtain the claim graph CG
3       $CG\leftarrow\text{cross\_claim\_extract}(Q||c_{i})$
       // Sample a subgraph $SG$ from $G$
4       $SG\leftarrow\text{sample\_graph}(G)$
       // Encode $SG$’s components
5       ${}{}SGE,CGE\leftarrow\text{\{\}},\text{\{\}}$} , {}
6       for $(E_{a},r_{ab},E_{b})\in SG$ do
7             $SGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
8       end for
      // Encode $CG$’s components
9       for $(E_{a},r_{ab},E_{b})\in CG$ do
10             $CGE\text{.append}(\text{enc}(E_{a}),\text{enc}(r_{ab}),\text{enc}(E_{b}))$
11       end for
      // Compute the alignment between encoded claims and all encoded relations
12       for $(h^{i}_{c},h^{j})\in\text{GAT}(SGE)\times\text{GAT}(CGE)$ do
13             $R^{ij}\leftarrow\cos(h^{i}_{c},h^{j})$
14
15       end for
      // Compute the embedding for all neighboring claims
16       $\bar{H}\leftarrow Rh$
       // Compute the final score using self-attention
17       $c\leftarrow\text{enc}(Q||C)$
18       $c_{\text{final}},H_{\text{final}}\leftarrow\text{SelfAttention}(c||\bar{H})$
       // Save the score of the current choice
19       $P\text{.append}(c_{\text{final}})$
20 end for